Please help me with the following ptoblem:
4^-1 - 2^-1

Answer by nerdybill(7384)

Remember: x^-1 can be rewritten as 1/x
.
4^-1 - 2^-1
= 1/4 - 1/2
= 1/4 - 2/4
= -1/4